3. a) Define crystal and unit cellย 

b) Explain de Broglie wave and the photoelectric effect.ย 

c) A wire of length 7 ๐‘š and radius 2.37 ๐‘š๐‘š is stretched by a load of 12370 ๐‘”๐‘š and

increased by 1.78% of the original length. Find the stress, strain, and Youngโ€™s

modulus of the material of the wire.


1
Expert's answer
2021-12-10T11:31:02-0500

a) Crystal is an arrangement of atoms that makes up a solid. Unit cell is the smallest arrangement of the crystal that acquires its properties.

b) De Broglie wave is a wave that would a body acquire if it were an elementary particle. The photoelectric effect is the emission of electrons when the light is shone on metals.

c) See the solution below:


"\epsilon=1.78\%/100\%=0.0178.\\\space\\
\sigma=\frac{F}{A}=\frac {mg}{\pi r^2}=6.87\text{ MPa}.\\\space\\

E=\frac{\sigma}{\epsilon}=386\text{ MPa}."